A pharmacist needs 70 liters of a 50% alcohol solution. She has availible a 30 % solution and an 80% solution. How many liters of each solution should she mix to obtain 70 liters of a 50% alcohol solution ?

Also from a take home exam

Answer by checkley77(12844)

70*.5=.30X+.80(70-X)
35=.30X+56-.80X
35-56=-.50X
-21-.5X
X=-21/-.5
X=42 LITERS OF THE 30% ALCOHOL IS NEEDED.
70-42=28 LITERS OF THE 50% ALCOHOL IS ALSO NEEDED.
PROOF:
70*.5=.30*42+.80*28
35=12.6+22.4
35=35